OpenStudy (anonymous):

find the vertex of the function given below ? y=x^2-2x+1

OpenStudy (anonymous):

i got i got 1,0

hero (hero):

The vertex of any parabola is \[x = -\frac{b}{2a}\] In other words, if you have a = 1 and b = -2, then \[x = -\frac{(-2)}{2(1)}\]

hero (hero):

So \[x = \frac{-(-2)}{2} = \frac{2}{2} = 1\]

hero (hero):

Now that we know x = 1, we insert that into the equation to find the y value.

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

we plug it in

hero (hero):

Sorry, I didn't see your first post.

OpenStudy (anonymous):

then we got zero